OpenStudy (anonymous):

Can somone help me with this word problem: Sand is falling of a converyor belt onto a conical pile at the rate of 15ft^3/min. The diameter of the base of the cone is approxiamtly twice the altitude. At what rate is the height of the pile changing when it is 10 feet high?

OpenStudy (anonymous):

You need to to relate the height to the diameter to the volume. The volume of a cone is: (1/3)pi r^2 h But you know that the DIAMETER is twice the HEIGHT. Or, the radius is 4 times as much. So: V=(1/3)pi (4h)^2*h V=(1/3)pi(16h^3) Differentiate it with respect to time. \[\frac{dV}{dt}=\frac{16 \pi}{3}3h^2\frac{dh}{dt}\] dV/dt=15ft^3/min h=10 Solve for dh/dt. \[15=16\pi(10)(\frac{dh}{dt}) \rightarrow \frac{dh}{dt}=\frac{15}{160 \pi}\]

OpenStudy (anonymous):

That what you got satellite?

OpenStudy (anonymous):

\[V=\frac{1}{3}\pi r^2 h =\frac{1}{3} \pi (4h)^2h=\frac{1}{3}\pi 16h^3\]\]

OpenStudy (anonymous):

yeah looks good to me.

OpenStudy (bahrom7893):

I confirm.

OpenStudy (anonymous):

I have the answer as being 0.048 ft/min

OpenStudy (anonymous):

\[V'=16 \pi h^2h'=15\]

OpenStudy (anonymous):

replace h by ten, solve for h' get just what you have. i am slow.

OpenStudy (anonymous):

oh hold on

OpenStudy (anonymous):

should be \[h'=\frac{15}{16 \pi ^2h}\] yes?

OpenStudy (anonymous):

so if \[h=10\] get \[h'=\frac{15}{1600 \pi}\]

OpenStudy (bahrom7893):

Sorry but malevolence was wrong.. I was wrong too..

OpenStudy (anonymous):

sorry i meant \[h'=\frac{15}{16 \pi h^2}\]

OpenStudy (anonymous):

replace h by ten get \[h'=\frac{15}{1600 \pi}\]

OpenStudy (anonymous):

i was able to get rid of one variable r. By saying that in this case d(diameter)=2h(height). And i know diameter=2 times radius, so 2r=2h which is r = h. SO the equation is: v=1/3*pi*h^3

OpenStudy (bahrom7893):

Okay guys look: V = (1/3)pie*r^2*h diameter is twice the height, therefore: d = 2h 2r = 2h r = h. So plug that in: V = (1/3)pie*h^2*h

OpenStudy (bahrom7893):

V = (pi/3 ) * h^3 dV/dt = pi * h^2 (dh/dt)

OpenStudy (anonymous):

Now the question i have is what does thhat equation look like when i differntiate it?

OpenStudy (bahrom7893):

15 = pi * 100 dh/dt dh/dt = 15/(100pi)

OpenStudy (anonymous):

yeah looks like we all blew it in two ways. \[r=h\] and so \[V=\frac{1}{3} \pi h^3\]

OpenStudy (anonymous):

Also i know dv/dt=15*ft^3/min

OpenStudy (bahrom7893):

I just posted the solution, dont u guys see it?

OpenStudy (anonymous):

He's right. Mine is wrong, I miswrote it :(

OpenStudy (anonymous):

\[V'=\pi h^2h'=15\] or \[h'=\frac{15}{\pi h^2}\]

OpenStudy (bahrom7893):

V = (1/3)pie*r^2*h diameter is twice the height, therefore: d = 2h 2r = 2h r = h. So plug that in: V = (1/3)pie*h^2*h V = (pi/3 ) * h^3 dV/dt = (3pi/3) * h^2 (dh/dt) dV/dt = pi * h^2 (dh/dt) dV/dt is 15, so 15 = pi * 10^2 dh/dt dh/dt = 15/(100pi)

OpenStudy (anonymous):

but how do i get rid of the cube part in 15ft^3/min

OpenStudy (bahrom7893):

what cube?

OpenStudy (anonymous):

that just means the units are cube units

OpenStudy (bahrom7893):

you mean the units? ft^3/min? ignore them...

OpenStudy (anonymous):

volume comes in the cube

OpenStudy (anonymous):

how come the answer does not have a cube in it?

OpenStudy (anonymous):

You have ft^3 for volume and you're dividing it by h^2 or ft^2 so ft^3/ft^2=ft

OpenStudy (bahrom7893):

the answer is the change in height.. change in height is in feet per minute.. height is not volume..

OpenStudy (anonymous):

malevolence nailed it. the units work themselves out

OpenStudy (bahrom7893):

Height is like length - feet Area - ft^2 Volume - ft^3

OpenStudy (bahrom7893):

can I get a medal too?

OpenStudy (anonymous):

exactly so if you take volume divided by area you get height.

OpenStudy (radar):

Medal awarded to bahroom

OpenStudy (bahrom7893):

lol tnx guys xD